OpenStudy (anonymous):

Determine whether the sequence is geometric. If it is, find the common ratio.

OpenStudy (anonymous):

a)\[-1,1,-1,1,...\] b) \[1, \frac{ 1 }{ 2 },\frac{ 1 }{ 3 },\frac{ 1 }{ 4 },...\]

OpenStudy (anonymous):

Use the rule outlined in my attachment hereto. Plug in the given values for each sequence to determine whether it a geometric sequence.

OpenStudy (anonymous):

I dont understand, can you work A and I will see what you do?

OpenStudy (campbell_st):

well to find the common ratio compare the terms if you have \[a_{1}, a_{2}, a_{3}, a_{4} .....\] look at the ratios, does \[\frac{a_{2}}{a_{1}} = \frac{a_{3}}{a_{2}} = \frac{a_{4}}{a_{3}}....\] so in you're 1st question does \[\frac{-1}{1} = \frac{1}{-1} = \frac{-1}{1}.....\] if the answer is yes or true... then you have a geometric sequence... hope it helps

OpenStudy (anonymous):

it is a geometric sequence, thats what I got. and the common ratio is 1?

OpenStudy (campbell_st):

not quite, it is geometric... but the common ratio is -1.... that why the sign of each term changes... so in the 2nd question does \[\frac{\frac{1}{2}}{1} = \frac{\frac{1}{3}}{\frac{1}{2}} = ...\] if it yes... its geometric... if now... its not...

OpenStudy (campbell_st):

oops no.... rather that now

OpenStudy (anonymous):

Alright, for B. It is not geometric. Is that correct?